OpenStudy (anonymous):

Write a recursive formula for the sequence 1, 3, 9, 27, 81, . . . a1 = 1

OpenStudy (jack1):

for all sequences, always check the difference between each successsive number to identify the pattern

OpenStudy (campbell_st):

\[\sum_{n=0}^{\infty} (3)^n\]

OpenStudy (anonymous):

\[a_{n} = r (a_{n - 1})\] where: r = common ratio so, to find the common ratio we do the following: \[r = \frac{ a _{n} }{ a _{n-1} } = \frac{ 3 }{ 1 } = 3\] \[r = \frac{ a _{n} }{ a _{n-1} } = \frac{ 9 }{ 3 } = 3\] \[r = \frac{ a _{n} }{ a _{n-1} } = \frac{ 27 }{ 9 } = 3\] So we can conclude that the common ratio is 3. \[a_{n} = 3(a_{n-1})\]

OpenStudy (anonymous):

so your recursive formula for the sequence would be: \[a_{n} = 3(a_{n - 1})\]